OpenStudy (anonymous):

Choose the correct description of the graph of the compound inequality x - 2 > -4 and 3x less than or greater to 15. A)A number line with an open circle on -2, shading to the left, and a closed circle on 5, shading to the right B)A number line with a closed circle on -2, shading to the left, and an open circle on 5, shading to the right C)A number line with an open circle on -2, a closed circle on 5, and shading in between D)A number line with a closed circle on -2, an open circle on 5, and shading in between

OpenStudy (anonymous):

@amorfide

OpenStudy (anonymous):

hey i need help on that @amorfide

OpenStudy (amorfide):

okay do you mean 3x less than or equal to 15

OpenStudy (anonymous):

sorry its this one x - 2 > -4 and 3x less than or greater to 15.

OpenStudy (amorfide):

i am going to repeat my question do you mean 3x less than or equal to 15

OpenStudy (amorfide):

\[3x \le 15\]

OpenStudy (amorfide):

do you mean that

OpenStudy (anonymous):

yes

OpenStudy (amorfide):

okay so

OpenStudy (amorfide):

first we need to solve for x x-2 > -4 solve for x

OpenStudy (anonymous):

gimmie a sec im tying to solve this

OpenStudy (anonymous):

its gunna be b

OpenStudy (amorfide):

no

OpenStudy (amorfide):

x-2>-4 x>-2 when we have greater than -2, it means it can not be equal to -2 but we want every value above -2 to put this on a number line to say we want greater than -2 but not equal to -2 we use an empty circle since the sign is greater than, then we want to shade to the right of -2 like this 

OpenStudy (amorfide):

now we have \[3x \le 15\] solve for x to get \[x \le 5\]

OpenStudy (amorfide):

when we want to draw equal to, on a number line, we do a circle that is shaded in and we want less than 5 also so we shade in the left side of 5